OpenStudy (anonymous):

What is the slope of a line perpendicular to the line with equation y = 6x + 2? A. -1/6 B. 6 C. -6 D. 2

OpenStudy (anonymous):

Someone?

OpenStudy (anonymous):

A.

OpenStudy (anonymous):

Negative recipicals when it's perpendicular.

OpenStudy (anonymous):

So its A

OpenStudy (anonymous):

Yes it is. *Reciprocals Lol

OpenStudy (anonymous):

Yes. y = mx + b perpendicular is y = -x/m + e

OpenStudy (anonymous):

Thanks I just don't get graphs....

OpenStudy (anonymous):

In time, with practice, you will.

OpenStudy (anonymous):

I have another question. Could you help me.

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

Question?

OpenStudy (anonymous):

Which of the following has a slope of –3 and a y-intercept of –4?

OpenStudy (anonymous):

y = mx + b has slope of m and y-intercept of b put these equations into this form and the answer becomes clear.

OpenStudy (anonymous):

A. y = 3x – 4 B. y = 3x + 4 C. y = –3x + 4 D. y = –3x – 4

OpenStudy (anonymous):

Those are the answers.

OpenStudy (anonymous):

Fine. You know how to find m and b y = (slope) x + (y-intercept)

OpenStudy (anonymous):

I believe so.

OpenStudy (anonymous):

Compare y = -3x -4 with the options

OpenStudy (anonymous):

@ColePhelps1 What's the answer????

OpenStudy (anonymous):

I got it "y=-3x-4"

OpenStudy (anonymous):

Right. Answer D